OpenStudy (anonymous):

Find the magnitude of the force needed to accelerate a 220g mass with a⃗ = -0.145m/s2 i^+0.430m/s2 j^.

OpenStudy (anonymous):

you cant find the magnitude of the acceleration vectors by using the pathagorean : (-.145)^2+(.430)^2=a^2 and solve for a and multiply this value by the mass in kilograms.

OpenStudy (anonymous):

can*

OpenStudy (anonymous):

\[\sqrt{(-.145)^{2}+(0.43)^{2}}=a(220*10^{-3})=\left| F \right| \]

OpenStudy (anonymous):

Can you just multiply the mass times the acceleration?

OpenStudy (anonymous):

yes, but you need the net acceleration. Take a look at the equation above.

OpenStudy (anonymous):

In that equation there are two variables, the a and the F?

OpenStudy (anonymous):

is the acceleration the answer in parenthesis?

OpenStudy (anonymous):

Sorry, I wrote the equation wrong. Once you solve for a, then you can multiply by the mass.

OpenStudy (anonymous):

the accelleration is computed with the pathagorean, and then you multiply the term in parentheses (mass converted to kilograms) to find the force. I'll rewrite the equation properly.

OpenStudy (anonymous):

I got .0999 N. is that correct?

OpenStudy (anonymous):

\[\sqrt{a _{x}^{2}+a _{y}^{2}}=a _{net},\left| F \right| =ma _{net}\]

OpenStudy (anonymous):

Not sure, let me calculate.

OpenStudy (anonymous):

that's it.

OpenStudy (anonymous):

and if i were to find the angle of this do i take the tangent of it?

OpenStudy (anonymous):

sorry not angle but direction

OpenStudy (anonymous):

you would take the arctangent of the two acceleration component vectors.

OpenStudy (anonymous):

You can figure out direction just by looking at the signs of the component vectors for the acceleration vector in relation to the cartesian coordinate grid.

OpenStudy (anonymous):

would the answer be 15.02 im not sure what the unit is

OpenStudy (anonymous):

no, I got the value 18.6 degrees.

OpenStudy (anonymous):

how do you convert that answer using dimensions of plane angle.

OpenStudy (anonymous):

I don't know what that means. What is plane angle?

OpenStudy (anonymous):

well the question is asking for the direction of the force \[\theta=\]

OpenStudy (anonymous):

then you would make the degree angle negative.

OpenStudy (anonymous):

Hey, sorry gave you the wrong angle earlier. I got -71.4 degrees

OpenStudy (anonymous):

\[\tan^{-1} (0.43/-0.145)=\theta \]

OpenStudy (anonymous):

i tried it and got the same answer to but it keeps saying its wrong, im not sure what exactly its asking for

OpenStudy (anonymous):

maybe it wants 108.6 degrees. since the vector should be in the second quadrant of the cartesian

OpenStudy (anonymous):

Yes that is the correct answer. How did you know it was suppose to be in the second quadrant

OpenStudy (anonymous):

just by looking at the vector signs. -+ is the second quadrant of the cartesian. -x, +y

OpenStudy (anonymous):

Awesome. I get it now. Thanks. :D

OpenStudy (anonymous):

no problem.